OpenStudy (amberlykhan):

How do you factor these: 1. 16b^2 - 4c^2 2. 100y^16 - 64x^16

jagr2713 (jagr2713):

Hey, 1. what can we factor out of the equation?

OpenStudy (amberlykhan):

I'm not sure...

jagr2713 (jagr2713):

Ok so what is the GCF of this equation or what number can go into both 4 and 16

OpenStudy (amberlykhan):

4

jagr2713 (jagr2713):

Great, any variables can be factored?

OpenStudy (amberlykhan):

b and c?

jagr2713 (jagr2713):

No, they arent like terms so just 4 so what is 4 divided by each number

OpenStudy (amberlykhan):

4 and 1

jagr2713 (jagr2713):

So what do we have left?

OpenStudy (amberlykhan):

4b^2 -c^2

jagr2713 (jagr2713):

you forgot the 4 so we have\[4(4b ^{2}-c ^{2})\]

jagr2713 (jagr2713):

Do you think we are done?

OpenStudy (amberlykhan):

no we have to factor out the expression in the parenthesis further

jagr2713 (jagr2713):

Whats it called? difference of ??

OpenStudy (amberlykhan):

ugh idk sorry

OpenStudy (whpalmer4):

if we have something \(x\) and we take \(x*x = x^2\) what have we done to \(x\)?

jagr2713 (jagr2713):

Its called difference of perfect squares

OpenStudy (amberlykhan):

ohhhhhhh

jagr2713 (jagr2713):

so a2-b2=(a+b)(a-b)

OpenStudy (amberlykhan):

ya

jagr2713 (jagr2713):

So whats 4 squared ?

OpenStudy (amberlykhan):

2

OpenStudy (whpalmer4):

\(4\) squared means \(4*4\)

jagr2713 (jagr2713):

How about C

OpenStudy (amberlykhan):

would it be (2b^2 - c) (2b^2 - c)

OpenStudy (amberlykhan):

i mean + c

OpenStudy (amberlykhan):

for the second

OpenStudy (amberlykhan):

and then add the 4 outside the parenthesis

jagr2713 (jagr2713):

Where id you get the 2b^2

jagr2713 (jagr2713):

You split the ^2 into two parenthesis

OpenStudy (amberlykhan):

bc 2b * 2b is 4b^2

jagr2713 (jagr2713):

Yes its 2b not 2b^2

OpenStudy (amberlykhan):

oh ya sorry thats what i meant

jagr2713 (jagr2713):

And + and -

jagr2713 (jagr2713):

So whats the final answer?

OpenStudy (amberlykhan):

4(2b -c) (2b + c)

jagr2713 (jagr2713):

Correct :D

jagr2713 (jagr2713):

Try the next one, tell me what you get

OpenStudy (amberlykhan):

ok

OpenStudy (amberlykhan):

I'm stuck at 4(24y^16 - 16x^16)

jagr2713 (jagr2713):

16 divdied by 2?

OpenStudy (amberlykhan):

8

OpenStudy (amberlykhan):

wouldnt it be -8

jagr2713 (jagr2713):

No because 16 is not negative only 16 is

OpenStudy (amberlykhan):

?

jagr2713 (jagr2713):

64 dvided by = 16 and its -16 but ^16 is positive

OpenStudy (amberlykhan):

but arent u dividing by -64?

jagr2713 (jagr2713):

Yes so we get -16x^16

OpenStudy (amberlykhan):

so what do we do after this 4(24y^16 - 16x^16)

OpenStudy (amberlykhan):

25

OpenStudy (amberlykhan):

4(25y^16 - 16x^16)

jagr2713 (jagr2713):

Remember difference of perfect squares ?

OpenStudy (amberlykhan):

ok lemme try

OpenStudy (amberlykhan):

i give up

jagr2713 (jagr2713):

Sorry i have finals tomorrow and i cant stay but just divided the exponents by 2 and you will get your answer and dont forget to sqrt 25 and 16

OpenStudy (whpalmer4):

@AmberlyKhan \[4(25y^{16} - 16x^{16})\]Let's just look at \(25y^{16}\)and see if we can recognize it as something that is a perfect square. what number times itself = 25? what power of \(y\) times itself = \(y^{16}\)? remember that when you multiply exponentials with the same base, you keep the base and add the exponents, so \(x^2*x^2 = x^{2+2} = x^4\)

OpenStudy (amberlykhan):

4(5y^8 - 4x^8) (5y^8+ 4x^8)

OpenStudy (whpalmer4):

yes, that is correct! You can check that is correct by multiplying it out: \[4(5y^8-4x^8)(5y^8+4x^8) = 4(25y^{16} + 20x^8y^8-20x^8y^8-16x^{16})\]\[=4(25y^{16}-16x^{16}) = 100y^{16}-64x^{16}\checkmark \]

OpenStudy (amberlykhan):

ok thx so much

jagr2713 (jagr2713):

Sorry for leaving, had to do some studying, but great job

jagr2713 (jagr2713):

\[ \huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\] if you want to know more about this site, Here is a tutorial about this site : http://openstudy.com/study#/updates/5525c0fce4b030ceb3f357e8